56) Geometry. A triangle has sides 3x + 7, 4x – 9, and 5x + 6. Find the expression that represents its perimeter.
Answer by checkley75(3666)

3X+7+4X-9+5X+6=12X+4=4(3X+1) IS THE PERIMETER.